Jerry is four times as old as Laura. In 20 years he will be twice as old as Laura. How old are Jerry and Laura?
Answer by ikleyn(52781)

.
Jerry is four times as old as Laura. In 20 years he will be twice as old as Laura. How old are Jerry and Laura?
~~~~~~~~~~~~~~~~~~~~~~~~

J = 4L.           (1)

In 20 years:

J+20 = 2(L+20).   (2)

Substitute (1) into (2). You ill get

4L + 20 = 2(L+20),  or

4L + 20 = 2L + 40,

4L - 2L = 40 - 20,

2L = 20  --->  L = 10.  Laura is 10 years now.

J = 4L = 4*10 = 40.  Jerry is 40 years now.

Answer. Laura is 10 years now.  Jerry is 40 years now.

Please check it yourself.

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

It contains a lot of other word problems and many other interesting and useful things.